From nch@roe.ac.uk Tue May 8 11:13:44 2007 Date: Fri, 4 May 2007 17:44:47 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Johann Bryant , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann , Ross Collins Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , Brian Walshe , John Taylor , Jim Emerson , Malcolm Stewart , Lorenzo Rimoldini , Mike Irwin , Mark Holliman , Peredur Williams , Stephen Warren Subject: WFAU VDFS Science Archive weekly project meeting minutes, 4 May 07 Minutes of WFAU VDFS Science Archive meeting: 4th May 2007 ------------------------------------------------------------------------- ------------------------------------------------------------------------- Present: NCH, PMW, RSC, MAR, RGM, ETWS, JMS, JB, LGR, MSH Apologies: JDT, BCW, JPE, AL, NJC DONM: 10am, Friday 11th May 2007 in the Plate Library Actions discharged this week: ----------------------------- ACTION: NCH to polish up CU19 for a few minor issues arising in the DR2 GPS run. Discharged with much help from RSC and advice from JB. Still needs tidying up for non-surveys, but this is a lower priority. ACTION: ETWS to email SJW as soon as some 06B data are available for perusal in the WSA. Discharged. ACTION: RSC to add new keywords into Multiframe, and check for any others. Discharged; much detailed discussion ensued concerning new keywords appearing in the files; these will be added in to the Multiframe schema and populated over the weekend. The general rule is: if in doubt, propagate all metadata into the relevant DB tables. Actions partly discharged but continuing: ----------------------------------------- The following from last time partly done but continue: o) Add in a default row for every detector appearing in every detection table (for schema consistency when querying merged sources and individual detections) - ACTION: RSC & NCH Continues Actions carried forward from 27/04/07 meeting: ---------------------------------------------- The following from last time continue: ACTION: NCH to discuss and review AstroWISE interfacing with MAR and JDT. ACTION: NCH to start some hard-nosed negotiations with the UKIDSS PI and CSS concerning the contents and timing of DR3. - CONTINUES; will start at next week's UKIDSS Survey Heads meeting. ACTION: AL to check on the exact wording of the survey release policy with reference to world release of proprietary survey data. - CONTINUES Specific points and new actions: -------------------------------- Project management: NCH and PMW noted VDUC next week. Attendees will be NCH, PMW; all others are invited to attend a buffet lunch to mingle with the committee. Location and time are the Management Building meeting room; kick-off 11am. ACTION: NCH to book a projector for the day ACTION: PMW to book buffet lunch and coffee for appropiate numbers. WFCAM & VISTA updates: JMS gave the team an update on Vista. Pointing/tracking are looking OK with TCS GUIs all integrated in nicely. Comments and issues arising from CASU fortnightly minutes: The team noted the minutes of the meeting of 26th April; the only comment was on the photometric calibration tweak (see Software below). Networking: JB, ETWS and MSH noted that UKLight is now up and running; 06B transfers will henceforth be done on this network. Some benchmarking will take place over the next few weeks as a part of normal transfer operations and possible optimisation will take place as necessary. Thanks to all involved for their tenacity in getting the link set up (EUCS, UKERNA, CASU and WFAU). WSA Operations: 06B transfer and ingest continues as normal. JB noted that he has brought the 06A non-survey situation under control with DBs prepared for most PIs; MAR has notified them of the availability. Hardware: New server still awaits suitable rackmount components. Looking further ahead, JB, ETWS and MSH noted that new storage and a new rack would likely be needed sooner rather than later. Software: Prompted by the communique from MJI concerning small photometric calibration tweaks as a function of field position, a debate ensued concerning how and when to apply this. The mod is easy enough to incorporate for ingests from now on, but NCH noted previous difficulties with transaction-heavy modification of catalogue calibrated attributes. The team agreed that such a modification is important and if applied from now on it must also be applied retrospectively. Major discussions concerning the practicalities were brought to a close by NCH who suggested that all concerned meet offline to brainstorm a general procedure for doing this (also with a view to astrometric recalibration at some point in the future). ACTION: Technical focus group to meet after next week's project meeting to sort out the general recalibration problem. MAR reported compiling usage stats for the WFAUs various archives for inclusion in grant renewal etc. RSC and ETWS have tweaked image metadata ingest to revert to CU3's original behaviour when given wacky WCS data. Survey Data Release: A lively discussion has occured concerning rationalisation of the GPS schema to expedite curation and querying. This will likely be drawn to a close next week when the GPS WG Survey Head visits Edinburgh for VDUC. Non-survey Data Release: JB noted that a large tranche of non-survey prepared DB releases has taken place over the last week which should satisfy a bundle of non-survey PIs. MAR noted updating of userInterface tables/logins etc for newly released nonSurvey databases. Astrogrid deployment: MSH has supplied some DSA usage stats to MAR for inclusion in current activity summaries, but also asked what kind of supplementary information might be required. Suggestions included breakdown by institute, user and dataset, as well as interface (ADQL or ConeSearch). Row counts and data volumes would also be very useful. Miscellaneous: The team congratulated LGR on his recent successful defense of his PhD thesis, and charged him with organising celebratory beers sometime next week.